OpenStudy (anonymous):

Solve the system of equations. 2x - 5y = 3 x - 3y = 1 (4, 1) (5, 2) (7, 2) (9, 3)

OpenStudy (sleepyhead314):

first isolate the x from the second equation by adding 3y to both sides of the second equation

OpenStudy (anonymous):

x=3y+1

OpenStudy (anonymous):

@sleepyhead314

OpenStudy (anonymous):

ummm are you there? or nah?

OpenStudy (sleepyhead314):

sorry, I was afk for a while yes x = 3y + 1 is correct :) then plug that into x into the first equation and solve for y

OpenStudy (anonymous):

How do i "plug it in"?

OpenStudy (sleepyhead314):

plug in (3y+1) for x into 2x - 5y = 3 like 2(3y+1) - 5y = 3

OpenStudy (anonymous):

I am stupid how would i solve that?

OpenStudy (sleepyhead314):

distribute the 2 then add like terms then subtract 2 from both sides :)

OpenStudy (anonymous):

i am soooo confused............

OpenStudy (sleepyhead314):

2(3y+1) - 5y = 3 distribute the 2 -> 6y + 2 - 5y = 3 do you see what I did there? :)

OpenStudy (anonymous):

yeahh

OpenStudy (sleepyhead314):

then add like terms 6y + 2 - 5y = 3 6y - 5y + 2 = 3 so y + 2 = 3 can you figure out what y equals from there? :P

OpenStudy (anonymous):

y=1

OpenStudy (sleepyhead314):

yay! correct :) well technically you can get the answer from there :P in the (x, y) format, when y=1 which answer choice do you think would match it?

OpenStudy (anonymous):

A!!!

OpenStudy (sleepyhead314):

Tah Dahhhh!!! xD that's correct ^_^

OpenStudy (anonymous):

! can you help me with one more prblem?

OpenStudy (sleepyhead314):

sorry, I've gotta go to sleep :/

OpenStudy (anonymous):

ok byye!